OpenStudy (anonymous):

Shannon and Macie are raking leaves to earn some money. Shannon can rake 2 lawns that are about the same size in 5 hours. Macie can rake 2 lawns that are about the same size in 6 hours. How long would it take both girls to work together to rake 2 lawns? Write and solve an equation for this situation. Explain how to set up the equation, using w = rt.

hero (hero):

Hint: \[\frac{(5)(6)}{5 + 6} = t\]

OpenStudy (anonymous):

rate=work/time If you need to know how long it takes to do 2 lawns, (that's a work of 2 lawns) then you need to set up the equations for w. w(shannon)+w(macie)=2 rate*time for shannon + rate*time for macie = 2 \[\frac{2}{5}t+\frac{2}{6}t=2\] Solving for t would get you the time it takes for both to do 2 lawns.